# Confronting compositional confusion through the characterisation of the sub-Neptune orbiting HD 77946

### v1

@@ 5 Discussion
Our analysis allowed us to estimate the mass of HD 77946 b with a precision better than 20%, and the radius with a precision better than 10%. Combining the modelled mass and radius, we find the planet density to be g cm. We place the mass and radius measurements of HD 77946 b into context and compare them to composition models in Figure 7, which shows an M-R plot for small planets ( and 20 ). Confirmed planets, with a mass and radius estimated to a precision better than 20% and 10% respectively, taken from the Extrasolar Planets Encyclopedia666http://www.exoplanet.eu are also shown. Furthermore, the coloured lines in the M-R plot show different compositions, taken from Zeng et al. (2019)777Models are available online at https://lweb.cfa.harvard.edu/~lzeng/planetmodels.html (dashed lines) and Lopez & Fortney (2014) (solid lines). We include both models due to the recent conclusion of Rogers et al. (2023) in which it is found that the H/He mass relations in Zeng et al. (2019) are not applicable to planets undergoing atmospheric evolution. There are a range of Zeng et al. (2019) models available dependent on temperature, often incorrectly selected based on the equilibrium temperature of the planet. These models assume constant specific entropy, so should instead be selected based on the temperature of the corresponding specific entropy at 100-bar level in the gas envelope. We include the Zeng et al. (2019) models at 1000K (based on HD 77946 b’s equilibrium temperature calculated in Section 5.1) in Figure 7 so that a direct comparison can be made to previous mass-radius studies in the literature, but note that the Lopez & Fortney (2014) models are preferred as the assumption of a constant planet age is made instead of constant specific entropy, which is more suitable for planets undergoing atmospheric mass-loss. As can be seen in Figure 7, HD 77946 b is consistent with having a sub-Neptune composition with a H/He atmosphere making up about 1% of its mass or with having a water-world composition with a steam atmosphere.
@@ 5.1 Interior Structure
The result of the internal structure modeling for HD 77946 b is shown in Figure 8. The planet shows a relatively small gaseous envelope for a Sub-Neptune of 0.02 M corresponding to less than 1% of its total mass. However, due to its very high equilibrium temperature (1248 K), the radius of the gas layer is 0.69 R, which corresponds to 1/4 of the HD 77946 b’s total radius. Compared to the Earth, thanks to a possibly thick water layer (), the iron core is proportionally smaller ( mass fraction in HD 77946 b compared to it being 1/3 of the Earth’s mass).
The amount of water on the planet is almost unconstrained () due to the presence of an atmosphere. As the planet is close to its star we could assume that the planet formed inside the ice-line, that is to say it is a water-poor planet. We tried to model this eventuality by constraining the maximum fraction of water to . In this case, the envelope mass is bigger ( M) in order to retrieve a similar planet density, as well as its radius (), with the same . Both scenarios are plausible, however we favor the water-rich environment because the water-poor environment requires a strong assumption. Although giving slightly different results, they are both able to describe a planet of this mass and this radius.
@@ 5.3 Prospects for atmospheric observations
We calculated the transmission spectroscopic metric (TSM) and the emission spectroscopic metric (ESM) outlined in Kempton et al. (2018), for follow-up atmospheric observations with JWST, but find that the TSM and ESM (56.81 and 5.63 respectively) do not meet the threshold presented (suggested threshold of TSM > 90 for 1.5 < R < 10 R, and ESM > 7.5). Even though HD 77946 b may not have a high enough TSM, or ESM, to be amongst the ‘best-in-class’ targets highlighted by Hord et al. (2023), it still has one of the highest TSMs for a sub-Neptune orbiting a hot star (K), as shown in Figure 9, so is still an interesting target for transmission spectroscopy. However, the brightness of the target, at K=7.6, would saturate many instrument modes on JWST, but we emphasise that whilst this target may not be suitable for JWST spectroscopy, a single transit should give a clear detection of the atmosphere if no clouds or hazes were present. We verified this with PandExo models with equilibrium chemistry and no clouds or hazes, and found that G140H (and potentially G235H) would be able to pick up water features in a H/He-dominated atmosphere without fully saturating the instrument, so studies of its atmosphere for other reasons than just spectrum investigations are also viable.
We also looked into the possibility of atmospheric He detection through transmission spectroscopy in the near-infrared. Scaling from Kirk et al. (2022), who constrain the peak helium absorption for WASP-52b (J = 10.5, T) to 0.31%, for HD 77946 (J = 7.9, T) we could expect about 0.07% precision with Keck/NIRSPEC on a similar signal for HD 77946. Thus, detection of a strong metastable He signal would be challenging but perhaps not impossible in multiple transits because of the bright star, given that Zhang et al. (2021) achieved better precision on 55 Cancri e. Still, some caution is warranted as our low log(R’) value indicates that HD 77946 does not exhibit strong chromospheric activity. Planets orbiting stars with log(R’) -5 are rarely detected to exhibit helium absorption (Bennett et al., 2023), and unfortunately, given the distance (pc) and this inactivity, Lyman- detection would be challenging due to interstellar absorption.
@@ 6 Conclusions
In this paper, we present the analysis of the TESS target HD 77946, a bright (V = 9.00) F5 star (M = 1.17 M, R = 1.31 R). Using the results of a HARPS-N GTO campaign, as well as 2 additional CHEOPS transits, alongside the 2 initial TESS campaigns (each lasting around 27 days), we find HD 77946 to host one small (2.7 R, 8.4 M) transiting sub-Neptune, with an orbital period of 6.53 days. The accuracy of our mass measurement for this planet allows us to constrain likely composition scenarios. The planet is located above the radius valley, indicating that it is a sub-Neptune, and since the radius valley is theorised to separate super-Earths and sub-Neptunes the composition of the planet can be theorised in this context. We furthermore compare the properties of HD 77946’s planet alongside those with similarly constrained radii and masses, and find that these suggest that HD 77946 b is a sub-Neptune with an iron-core fraction smaller than that of the Earth and an H/He atmosphere that is about 1% of the total mass.

### v2

@@ 5 Discussion
Our analysis allowed us to estimate the mass of HD 77946 b with a precision better than 20%, and the radius with a precision better than 10%. Combining the modelled mass and radius, we find the planet density to be g cm-3. We place the mass and radius measurements of HD 77946 b into context and compare them to composition models in Figure 7, which shows an M-R plot for small planets ( and 20 ). Confirmed planets, with a mass and radius estimated to a precision better than 20% and 10% respectively, taken from the Extrasolar Planets Encyclopedia666http://www.exoplanet.eu are also shown. Furthermore, the coloured lines in the M-R plot show different compositions, taken from Zeng et al. (2019)777Models are available online at https://lweb.cfa.harvard.edu/~lzeng/planetmodels.html (dashed lines) and Lopez & Fortney (2014) (solid lines). We include both models due to the recent conclusion of Rogers et al. (2023) in which it is found that the H/He mass relations in Zeng et al. (2019) are not applicable to planets undergoing atmospheric evolution. There are a range of Zeng et al. (2019) models available dependent on temperature, often incorrectly selected based on the equilibrium temperature of the planet. These models assume constant specific entropy, so should instead be selected based on the temperature of the corresponding specific entropy at 100-bar level in the gas envelope. We include the Zeng et al. (2019) models at 1000K (based on HD 77946 b’s equilibrium temperature calculated in Section 5.1) in Figure 7 so that a direct comparison can be made to previous mass-radius studies in the literature, but note that the Lopez & Fortney (2014) models are preferred as the assumption of a constant planet age is made instead of constant specific entropy, which is more suitable for planets undergoing atmospheric mass-loss. As can be seen in Figure 7, HD 77946 b is consistent with having a sub-Neptune composition with a H/He atmosphere making up about 1% of its mass or with having a water-world composition with a steam atmosphere.
@@ 5.1 Interior Structure
The result of the internal structure modeling for HD 77946 b is shown in Figure 8. The planet shows a relatively small gaseous envelope for a Sub-Neptune of 0.02 M⊕ corresponding to less than 1% of its total mass. However, due to its very high equilibrium temperature (1248 K), the radius of the gas layer is 0.69 R⊕, which corresponds to 1/4 of the HD 77946 b’s total radius. Compared to the Earth, thanks to a possibly thick water layer (), the iron core is proportionally smaller ( mass fraction in HD 77946 b compared to it being 1/3 of the Earth’s mass).
The amount of water on the planet is almost unconstrained () due to the presence of an atmosphere. As the planet is close to its star we could assume that the planet formed inside the ice-line, that is to say it is a water-poor planet. We tried to model this eventuality by constraining the maximum fraction of water to . In this case, the envelope mass is bigger ( M⊕) in order to retrieve a similar planet density, as well as its radius (), with the same . Both scenarios are plausible, however we favor the water-rich environment because the water-poor environment requires a strong assumption. Although giving slightly different results, they are both able to describe a planet of this mass and this radius.
@@ 5.3 Prospects for atmospheric observations
We calculated the transmission spectroscopic metric (TSM) and the emission spectroscopic metric (ESM) outlined in Kempton et al. (2018), for follow-up atmospheric observations with JWST, but find that the TSM and ESM (56.81 and 5.63 respectively) do not meet the threshold presented (suggested threshold of TSM > 90 for 1.5 < Rb < 10 R⊕, and ESM > 7.5). Even though HD 77946 b may not have a high enough TSM, or ESM, to be amongst the ‘best-in-class’ targets highlighted by Hord et al. (2023), it still has one of the highest TSMs for a sub-Neptune orbiting a hot star (K), as shown in Figure 9, so is still an interesting target for transmission spectroscopy. However, the brightness of the target, at K=7.6, would saturate many instrument modes on JWST, but we emphasise that whilst this target may not be suitable for JWST spectroscopy, a single transit should give a clear detection of the atmosphere if no clouds or hazes were present. We verified this with PandExo models with equilibrium chemistry and no clouds or hazes, and found that G140H (and potentially G235H) would be able to pick up water features in a H/He-dominated atmosphere without fully saturating the instrument, so studies of its atmosphere for other reasons than just spectrum investigations are also viable.
We also looked into the possibility of atmospheric He detection through transmission spectroscopy in the near-infrared. Scaling from Kirk et al. (2022), who constrain the peak helium absorption for WASP-52b (J = 10.5, T) to 0.31%, for HD 77946 (J = 7.9, T) we could expect about 0.07% precision with Keck/NIRSPEC on a similar signal for HD 77946. Thus, detection of a strong metastable He signal would be challenging but perhaps not impossible in multiple transits because of the bright star, given that Zhang et al. (2021) achieved better precision on 55 Cancri e. Still, some caution is warranted as our low log(R’HK) value indicates that HD 77946 does not exhibit strong chromospheric activity. Planets orbiting stars with log(R’HK) -5 are rarely detected to exhibit helium absorption (Bennett et al., 2023), and unfortunately, given the distance (pc) and this inactivity, Lyman- detection would be challenging due to interstellar absorption.
@@ 6 Conclusions
In this paper, we present the analysis of the TESS target HD 77946, a bright (V = 9.00) F5 star (M⋆ = 1.17 M⊙, R⋆ = 1.31 R⊙). Using the results of a HARPS-N GTO campaign, as well as 2 additional CHEOPS transits, alongside the 2 initial TESS campaigns (each lasting around 27 days), we find HD 77946 to host one small (2.7 R⊕, 8.4 M⊕) transiting sub-Neptune, with an orbital period of 6.53 days. The accuracy of our mass measurement for this planet allows us to constrain likely composition scenarios. The planet is located above the radius valley, indicating that it is a sub-Neptune, and since the radius valley is theorised to separate super-Earths and sub-Neptunes the composition of the planet can be theorised in this context. We furthermore compare the properties of HD 77946’s planet alongside those with similarly constrained radii and masses, and find that these suggest that HD 77946 b is a sub-Neptune with an iron-core fraction smaller than that of the Earth and an H/He atmosphere that is about 1% of the total mass.
@@ Acknowledgements
This research has made use of data obtained from or tools provided by the portal exoplanet.eu of The Extrasolar Planets Encyclopaedia.